Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 1,000 hours each month to produce 2,000 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 41,400 $ 20.70
Direct labor $ 8,000 4.00
Variable manufacturing overhead (based on direct labor-hours) $ 3,400 1.70
$ 26.40

During August, the factory worked only 1,050 direct labor-hours and produced 2,400 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (7,500 yards) $ 48,000 $ 20.00
Direct labor $ 10,080 4.20
Variable manufacturing overhead $ 5,040 2.10
$ 26.30

At standard, each set of covers should require 3.0 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

1. Materials price variance
Materials quantity variance
2. Labor rate variance
Labor efficiency variance
3. Variable overhead rate variance
Variable overhead efficiency variance
Answer #1
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($6.90/Yards - $6.40/Yards ) × 7500Yards = $3750 F
* Actual price per Yard= 48000/7500= $6.4 per Yards
Std Price per Yards= $20.70/3 yeard= $6.90/ Yard
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((3 Yds   X 2400 Unit)-7500yds )X $6.90 = $2070 UF
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($8/hour - $9.60/hour) ×1050 Hours= $1680U F
*Actual Rate per Hour= 10080/ 1050=$9.60 per Hour
Std Rate per Hour= 8000/1000= $8per Hour
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × Std Hourly Rate
= (0.5 Hour × 2400 Unit - 1050) × $8 = $1200 F
Std Hour required per set of Cover= 1000/2000=0.5 hour/ set of Cover
Calculation of Variable OH rate variance
= (Standard Variable OH rate per hour - actual Variable OH rate per hour) × Actual hours used
= ($3.40/hour - $4.8/hour) ×1050 Hours= $1470 UF
*Actual Rate = $5040/1050=$4.80 / Hour
Std Rate per Hour= 3400/1000=$3.40/ Hour
Calculation of Vaaariable OH efficiency variance:
= (standard hours required for actual production - actual hours used) × Std Hourly Rate
= (0.5 Hour × 2400 Unit - 1050) × $3.40 = $510 F